OpenStudy (anonymous):

The table shows the number of candies packed by Machine C. The equation shows the number of candies packed by Machine D. In both representations, x is a measure of the number of minutes and y is a measure of the number of candies packed. Machine C Candy Packing x (minutes) y (candies) 2 150 4 300 6 450 8 600 Machine D: y = 130x How many more candies could machine D pack than machine C in 11 minutes? 65 130 450 605

OpenStudy (anonymous):

First, let's create an equation to represent the table. y = 75x

OpenStudy (anonymous):

Then, plug in 11 for x in both equations

OpenStudy (anonymous):

y = 75(11) y = 130(11) Solve the equations above, and let me know what you get. :-)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

I got 130 @bluediamond1999

OpenStudy (anonymous):

Not quite. y = 75(11) y = 825 y = 130(11) y = 1430 What is 1430 - 825?

OpenStudy (anonymous):

The answer is 605

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

No problem! Happy to help!